OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one third x - 6 y = -one third x - 14 y = 3x - 6 y = 3x - 14 @TheSmartOne

OpenStudy (anonymous):

First rearrange the equation to slope-intercept form. 3x - y = -3 Subtract 3x to both sides: -y = -3x - 3 Multiply -1 to both sides: y = 3x + 3 Now it's in the form of y = mx + b, where m is the slope. So what's the slope here? @autinregister